Suppose the dollar/franc exchange rate equals S0.50 parity (PPP) theory, what will happen to the dollars exchange value under each of the following per frane. According to the puachasing power circumstances? 1. (5 points) The U.S. price level increases by 10 percent, and the peice level in Switzerland increase by 5 percent. 2. (5 points) The U.S. price level increases by 10 percent, and the price level in Switzerland decreases by 5 percent 3. (5 points) The U.S. price level decreases by 10 percent, and the price level in decreases by 5 percent. 4. (S points) The U.S. price level decreases by 10 percent, and the price level in Switaerland increases by 5 percent 第4页
Answer #1

As we know that price level varies from one country to other country, in order to establish exchange rate equilibrium between different countries, we use the purchasing power parity (PPP) theory.

1. The U.S. price level increase by 10 percent, and the price level in switzerland increase by 5 percent, the new dollar's exchange value would be :

\bg_white The\ current\ exchange\ rate\ \times \frac{1+change\ in\ price\ index\ of\ U.S. after\ inflation}{1+change\ in\ price\ index\ of\ Switzerland\ after\ inflation}

The\ new\ dollar's\ exchange\ value= \$ 0.50\times \frac{1+.10}{1+.05}= \$ 0.52

Therefore, the new dollar's exchange value is $ 0.52/ franc.

2. The U.S. price level increase by 10 percent, and the price level in switzerland decrease by 5 percent, the new dollar's exchange value would be :

\bg_white The\ current\ exchange\ rate\ \times \frac{1+change\ in\ price\ index\ of\ U.S. after\ inflation}{1+change\ in\ price\ index\ of\ Switzerland\ after\ inflation}

The\ new\ dollar's\ exchange\ value= \$ 0.50\times \frac{1+.10}{1-.05}= \$ 0.58

Therefore, the new dollar's exchange value is $ 0.58/ franc.

3. The U.S. price level decrease by 10 percent, and the price level in switzerland decrease by 5 percent, the new dollar's exchange value would be :

\bg_white The\ current\ exchange\ rate\ \times \frac{1+change\ in\ price\ index\ of\ U.S. after\ inflation}{1+change\ in\ price\ index\ of\ Switzerland\ after\ inflation}

The\ new\ dollar's\ exchange\ value= \$ 0.50\times \frac{1-.10}{1-.05}= \$ 0.47

Therefore, the new dollar's exchange value is $ 0.47/ franc.

4. The U.S. price level decrease by 10 percent, and the price level in switzerland increase by 5 percent, the new dollar's exchange value would be :

\bg_white The\ current\ exchange\ rate\ \times \frac{1+change\ in\ price\ index\ of\ U.S. after\ inflation}{1+change\ in\ price\ index\ of\ Switzerland\ after\ inflation}

The\ new\ dollar's\ exchange\ value= \$ 0.50\times \frac{1-.10}{1+.05}= \$ 0.43

Therefore, the new dollar's exchange value is $ 0.43/ franc.